OpenStudy (anonymous):

In a plane, line b is perpendicular to line f, line f is perpendicular to line g, and line h is parallel to line f. Which of the following must be true? hg g || f b || h fh

OpenStudy (anonymous):

In a plane, line b is perpendicular to line f, line f is perpendicular to line g, and line h is parallel to line f. Which of the following must be true? hg g || f b || h fh

OpenStudy (anonymous):

Which is the equation of the line with slope 3 that contains point (-1, 5)? y – 5 = 3(x + 1) y – 1 = 3(x + 5) y – 1 = 3(x – 5) y – 5 = 3(x - 1)